a father gives his son certain amount of pocket money every week son saves 30% of amount and spends the rest with spending on saturday and sunday being twice than what is spent on oher days of the week if total saving in the week are 420 how much is spent in each saturday




Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

He gets x amount each week.
He saves 0.3x,
He spends 0.7x, which its 5 days a week the same and 2 days a week twice as much.
if what he spends a day is s then he spends 9s a week 5 days *1s+2 days*2s
so s=0.7x/9
0.3x=420, so x=1400
0.7x=980
s= spent per day and 2s each weekend day.
9s=980
s=108 8/9 a day
2s=217 7/9 a day on weekends
Total spending is 980
Total saving its 420
Total given is 1400 per week
-answer is 217 7/9






Answer by greenestamps(13327)



He saves 30%, so he spends 70%.

The 30% he saves is 420, so the amount he gets from his father each week is 420/.30 = 1400.

He spends 70% of the 1400, which is .70(1400)=980.

He spends x amount each day Monday through Friday and 2x amount each Saturday and Sunday, so the total amount he spends each week is 9x. The fraction of his weekly spending that occurs on Saturday is 2x/9x = 2/9.

(2/9)*980 = 1960/9 = 217 7/9.

ANSWER: 217 7/9

Probably not a reasonable answer; so perhaps you haven't shown the right numbers, or perhaps the problem is faulty....